### v1

@@ Abstract
Penetration testing, a critical component of cybersecurity, typically requires extensive time and effort to find vulnerabilities. Beginners in this field often benefit from collaborative approaches with the community or experts. To address this, we develop CIPHER (Cybersecurity Intelligent Penetration-testing Helper for Ethical Researchers), a large language model specifically trained to assist in penetration testing tasks. Unlike software development, penetration testing involves domain-specific knowledge that is not widely documented or easily accessible, necessitating a specialized training approach for AI language models. We trained CIPHER using over 300 high-quality write-ups of vulnerable machines, hacking techniques, and documentation of open-source penetration testing tools. Additionally, we introduced the Findings, Action, Reasoning, and Results (FARR) Flow augmentation, a novel method to augment penetration testing write-ups to establish a fully automated pentesting simulation benchmark tailored for large language models. This approach fills a significant gap in traditional cybersecurity Q&A benchmarks and provides a realistic and rigorous standard for evaluating AI’s technical knowledge, reasoning capabilities, and practical utility in dynamic penetration testing scenarios. In our assessments, CIPHER achieved the best overall performance in providing accurate suggestion responses compared to other open-source penetration testing models of similar size and even larger state-of-the-art models like Llama 3 70B and Qwen1.5 72B Chat, particularly on insane difficulty machine setups. This demonstrates that the current capabilities of general large language models (LLMs) are insufficient for effectively guiding users through the penetration testing process. We also discuss the potential for improvement through scaling and the development of better benchmarks using FARR Flow augmentation results. Our benchmark will be released publicly at https://github.com/ibndias/CIPHER.
@@
penetration testing; large language model; pentesting LLM, AI penetration testing assistant, domain specific LLM; LLM evaluation; vulnerabillity detection.
@@ 1 Introduction
Penetration testing, a fundamental aspect of cybersecurity, involves probing computer systems Denis et al. (2016), networks Jayasuryapal et al. (2021), or web applications Altulaihan et al. (2023) to identify security vulnerabilities that malicious actors could exploit. This proactive approach is critical for uncovering potential weaknesses before they can be manipulated, thus safeguarding sensitive data and maintaining the integrity of digital systems. Despite its importance, penetration testing is notoriously time-consuming and labor-intensive Clintswood et al. (2023), often requiring expertise and extensive collaboration. Traditional methods primarily rely on manual testing, which can be inefficient and may not scale well with modern IT infrastructures’ growing complexity and rapid evolution. As cyber threats continue to advance in sophistication, there is an increasing need for more efficient, comprehensive, and adaptable penetration testing methodologies.
Various approaches have been proposed to streamline the penetration testing process. Automated tools and frameworks, such as Metasploit Kennedy et al. (2011) and OpenVAS GmbH (2024), offer some level of automation and guidelines, yet they often require significant expertise to operate effectively and may not fully replace the nuanced insights provided by experienced testers. These tools can assist in discovering and exploiting vulnerabilities but are limited in their ability to provide comprehensive, adaptive, and intelligent responses to the dynamic nature of penetration testing tasks.
@@
The application of large language models (LLMs) in penetration testing represents a cutting-edge area of research with promising results. Advanced models such as GPT-4 Achiam et al. (2023) have demonstrated significant potential to automate and improve various penetration testing tasks, including Linux privilege escalation and the identification of file-based vulnerabilities Happe et al. (2024). However, implementing a straightforward LLM pipeline presents challenges, particularly in maintaining long-term memory for generating consistent and accurate commands throughout extended testing sessions Deng et al. (2023). Recent innovations like PentestGPT Deng et al. (2023) and AutoAttacker Xu et al. (2024) have addressed these limitations. These systems leverage existing LLMs and open-source models, incorporating specialized frameworks to emulate human penetration testers’ workflow and decision-making processes. While these advancements mark significant progress, it is important to note that these general-purpose models are not explicitly fine-tuned for the nuances of penetration testing. This lack of domain-specific optimization can potentially limit their effectiveness in handling complex, context-dependent scenarios often encountered in real-world penetration testing environments.
@@
Recognizing these limitations, we developed CIPHER (Cybersecurity Intelligent Penetration-testing Helper for Ethical Researchers). CIPHER is a large language model specifically trained and fine-tuned on a comprehensive penetration testing dataset. Leveraging advancements in generative AI, CIPHER aims to automate and enhance the penetration testing process, making it more efficient and accessible, especially for individuals with limited penetration testing knowledge. By fine-tuning the model with over 300 write-ups of vulnerable machines, hacking techniques, and documentation of open-source penetration tools, CIPHER provides detailed explanations and mimics the reasoning abilities of an expert penetration tester.
@@
Building on this foundation, we introduced the Findings, Action, Reasoning, and Results augmentation (FARR) Flow. The FARR Flow is a novel method that augments write-up documents for more effective training. It enhances the model’s ability to understand and replicate penetration testing processes and establishes a fully automated pentesting simulation benchmark tailored for large language models. This benchmark provides a realistic and rigorous standard for evaluating the technical knowledge, reasoning capabilities, and practical utility of AI in penetration testing scenarios.
@@
  • The methodology of CIPHER development, where we utilize write-ups into beginner and expert conversation to train a large language model tailored for practical penetration testing assistance.
  • Novel augmentation method to transform write-ups into compact information based on the Findings, Action, Reasoning, and Results (FARR) Flow augmentation method.
  • The development and evaluation of an automated pentesting simulation benchmark utilizing FARR Flow augmentation, providing a rigorous and realistic standard for assessing LLM performance in practical penetration testing guidance accuracy.
  • An extensive evaluation demonstrating CIPHER’s effectiveness and improvement over other penetration testing models.
These contributions bridge the gap between AI advancements and practical cybersecurity applications, offering a new penetration testing domain-specific language model and benchmark for realistic penetration testing scenarios.
@@ 2.1.1 Large Language Models
Transformers: Introduced by Vaswani et al. Vaswani et al. (2017) in "Attention is All You Need" (2017), Transformers have reshaped NLP by employing self-attention mechanisms to process entire sequences concurrently. They excel at capturing long-range dependencies, outperforming sequential models like RNNs. Transformers comprise self-attention layers, feed-forward networks, and positional encoding to maintain word order. Variants such as BERT and GPT have set benchmarks in NLP and extended into computer vision tasks through Vision Transformers (ViTs).
Open-Source LLMs: Open-source LLMs like the Llama family by Meta AI Touvron et al. (2023), Mistral Jiang et al. (2023), Yi 1.5 01.AI et al. (2024), and Qwen 1.5 Bai et al. (2023) democratize access to powerful language models. Evolving from Llama 1 to Llama 3, these models increase token size and context length to enhance language understanding. Innovations like the mixture of experts in Mixtral Jiang et al. (2024) and unique training methods in Yi 1.5 push the boundaries further. Qwen 1.5 excels in multilingual contexts, incorporates agentic frameworks Qwen Team (2024a), releasing smaller and powerful models Qwen Team (2024b), making them versatile tools Touvron et al. (2023); Mann et al. (2020).
@@ 2.1.2 Applications and Challenges in Cybersecurity
In cybersecurity, Large Language Models (LLMs) such as GPT-4Achiam et al. (2023), PentestGPTDeng et al. (2023), and HackerGPTMetta et al. (2024) have been deployed to assist with penetration testing and other security tasks. These models can analyze and synthesize large volumes of information, aiding in identifying vulnerabilities. However, using such advanced models poses significant challenges. The high cost and privacy risks associated with proprietary models like GPT-4, which handle sensitive vulnerability data, are significant concerns. Moreover, general-purpose LLMs often lack the specialized knowledge required for effective penetration testing.
@@ 2.2 Existing Tools and Methodologies for Penetration Testing
2.2.1 Traditional and Automated Tools
Penetration testing is a crucial component of cybersecurity, involving simulated cyberattacks to uncover system vulnerabilities. Traditional tools such as Metasploit Holik et al. (2014) and OpenVAS Aksu et al. (2019) offer robust vulnerability discovery and exploitation frameworks. These tools, however, heavily rely on user expertise to interpret results and often require manual intervention to execute complex attack scenarios.
Automated tools have emerged to streamline the penetration testing process by automating vulnerability scanning and initial exploitation attempts. Examples include Nessus Kumar (2014) and Burp Suite Kim (2020), which use predefined algorithms to detect and exploit common vulnerabilities. While these tools enhance efficiency, they may overlook subtle vulnerabilities that require human intuition and context to identify.
Combining traditional and automated tools optimizes penetration testing effectiveness. Traditional methods leverage human expertise to craft targeted attacks and interpret results, while automation accelerates routine tasks and broadens the scope of vulnerability detection. This hybrid approach ensures comprehensive security assessments addressing known vulnerabilities and emerging threats.
@@ 2.2.2 Advancements in AI for Penetration Testing
Recent advancements have seen the development of AI-driven tools designed to enhance penetration testing. Research efforts like PentestGPT Deng et al. (2023) and AutoAttacker Xu et al. (2024) leverage LLMs for various penetration testing tasks. Despite their potential, these tools face several limitations:
@@
  • Inaccurate Operations and Commands: LLMs may generate inaccurate commands or misinterpret outputs, leading to ineffective testing Fang et al. (2024a)Fang et al. (2024b).
@@ 2.2.3 Leveraging LLMs for Penetration Testing
On the other hand, general LLMs Achiam et al. (2023) used as penetration testing agents Deng et al. (2023)Fang et al. (2024a)Fang et al. (2024b) provide greater versatility and scalability. These models can perform a broader range of tasks beyond penetration testing and can be scaled easily by integrating additional modules or upgrading to more advanced general models. However, they often require complex setups involving various frameworks, modules, and pipelines, which can be resource-intensive and challenging to manage. The increased complexity can introduce vulnerabilities and make troubleshooting more difficult. Without significant adjustments and configurations, general LLMs may not achieve the same performance and specialization as native penetration testing models.
Table 1: Comparison of Current Penetration Testing LLMs. Notes (: available; : not available)
Name Native LLM Foothold Privilege Escalation Multiple Tools
@@
LLM Agents Can Autonomously Fang et al. (2024a)Fang et al. (2024b)
LLM as Hackers Happe et al. (2024)
@@
Given these considerations, we develop a native LLM tailored for penetration testing. This approach ensures our model is optimized for the unique requirements of penetration testing tasks, thereby achieving higher accuracy and efficiency. A specialized model integrates domain-specific knowledge and tools more effectively, simplifying deployment and reducing the need for extensive customization. Additionally, handling sensitive data locally enhances security and minimizes the risk of exposure to vulnerable information.
@@ 2.3.2 Evaluation and Benchmarking Challenges
Current evaluation frameworks for cybersecurity models, such as PurpleLlama CyberSecEval with MITRE ATT&CK, primarily focus on pre-defined scenarios and known attack patterns Strom et al. (2018). These do not adequately measure a model’s ability to discover novel vulnerabilities or adapt to dynamic situations encountered in real-world penetration testing. Previous works like PentestGPT and AutoAttacker lack reproducible benchmarks, relying on manual testing and human scoring, which introduces bias and limits reproducibility Deng et al. (2023); Xu et al. (2024). Developing standardized, automated benchmarks that reflect real-world conditions remains a significant challenge for advancing the evaluation of cybersecurity LLMs.
@@ 2.3.3 Reproducible Benchmark for Penetration Testing Scenario
To address these challenges, we introduce FARR Flow augmentation method, which automates the generation of penetration testing scenarios using high-quality write-ups. Unlike other existing penetration testing benchmarks for LLM, FARR Flow Evaluation covers a wide range of tool usage capabilities, from reconnaissance to exploitation tools. FARR Flow Evaluation results are also easily reproducible due to the scoring automation by using judge LLM. We also release our code on GitHub. This approach provides a dynamic and reproducible benchmark, assessing the LLM’s ability to reason and act like an expert penetration tester.
Table 2: Comparison of Penetration Testing Evaluation Approaches. Notes (: yes; : not provided; )
@@
LLM Can Autonomously Fang et al. (2024a)Fang et al. (2024b)
LLM as Hackers Happe et al. (2024)
@@
As shown in Table 2, our research demonstrates that FARR Flow evaluation fills the existing gap in current penetration testing benchmark for LLM that is easily reproducible, open-source and covers wide range of tools.

### v2

@@ Abstract
Penetration testing, a critical component of cybersecurity, typically requires extensive time and effort to find vulnerabilities. Beginners in this field often benefit from collaborative approaches with the community or experts. To address this, we develop Cybersecurity Intelligent Penetration-testing Helper for Ethical Researchers (CIPHER), a large language model specifically trained to assist in penetration testing tasks as a chatbot. Unlike software development, penetration testing involves domain-specific knowledge that is not widely documented or easily accessible, necessitating a specialized training approach for AI language models. CIPHER was trained using over 300 high-quality write-ups of vulnerable machines, hacking techniques, and documentation of open-source penetration testing tools augmented in an expert response structure. Additionally, we introduced the Findings, Action, Reasoning, and Results (FARR) Flow augmentation, a novel method to augment penetration testing write-ups to establish a fully automated pentesting simulation benchmark tailored for large language models. This approach fills a significant gap in traditional cybersecurity Q&A benchmarks and provides a realistic and rigorous standard for evaluating LLM’s technical knowledge, reasoning capabilities, and practical utility in dynamic penetration testing scenarios. In our assessments, CIPHER achieved the best overall performance in providing accurate suggestion responses compared to other open-source penetration testing models of similar size and even larger state-of-the-art models like Llama 3 70B and Qwen1.5 72B Chat, particularly on insane difficulty machine setups. This demonstrates that the current capabilities of general large language models (LLMs) are insufficient for effectively guiding users through the penetration testing process. We also discuss the potential for improvement through scaling and the development of better benchmarks using FARR Flow augmentation results.
@@
penetration testing; large language model; pentesting LLM; AI penetration testing assistant; domain specific LLM; LLM evaluation; vulnerabillity detection
@@ 1 Introduction
Penetration testing, a fundamental aspect of cybersecurity, involves probing computer systems Denis et al. (2016), networks Jayasuryapal et al. (2021), or web applications Altulaihan et al. (2023) to identify vulnerabilities that malicious actors could exploit. This proactive approach helps to uncover weaknesses before they can be manipulated, safeguarding sensitive data and maintaining system integrity. Despite its critical importance, penetration testing presents significant challenges, particularly for beginners. These challenges include understanding the complexity of modern IT infrastructures, mastering a wide range of tools, and gaining the necessary technical expertise to identify and exploit vulnerabilities effectively. Traditional penetration testing methods, which primarily rely on manual techniques, can be inefficient and labor-intensive Clintswood et al. (2023), making them especially difficult for newcomers. Moreover, scaling these methods to meet the complexity and rapid evolution of modern IT systems further complicates the process. As cyber threats continue to grow in sophistication, there is a growing need for penetration testing methodologies that are not only efficient and comprehensive but also accessible to beginners, helping them overcome the steep learning curve.
@@
Automated tools and frameworks, such as Metasploit Kennedy et al. (2011) and OpenVAS GmbH (2024), are designed to assist beginners by streamlining the penetration testing process. However, even with automation, users still need a certain level of expertise to effectively utilize these tools in dynamic pentesting situations, especially when interpreting and acting on the collected information. Meanwhile, the application of large language models (LLMs) in penetration testing represents a cutting-edge area of research with promising results. Advanced models such as GPT-4 Achiam et al. (2023) have demonstrated significant potential to automate and improve various penetration testing tasks, including Linux privilege escalation and the identification of file-based vulnerabilities Happe et al. (2024). However, implementing a straightforward LLM pipeline presents challenges, particularly in maintaining long-term memory for generating consistent and accurate commands throughout extended testing sessions Deng et al. (2023). Recent innovations like PentestGPT Deng et al. (2023) and AutoAttacker Xu et al. (2024) have addressed these limitations. These systems leverage existing LLMs and open-source models, incorporating specialized frameworks to emulate human penetration testers’ workflow and decision-making processes. While these advancements mark significant progress, it is important to note that these general-purpose models are not explicitly fine-tuned for the nuances of penetration testing. This lack of domain-specific optimization can potentially limit their effectiveness in handling complex, context-dependent scenarios often encountered in real-world penetration testing environments.
@@
Recognizing these limitations, we developed Cybersecurity Intelligent Penetration-testing Helper for Ethical Researchers (CIPHER) as a chatbot assistant specifically designed to support penetration testing tasks. CIPHER is a large language model fine-tuned on a specialized penetration testing dataset to assist ethical hackers by providing insights and recommendations. Trained on over 300 write-ups of vulnerable machines, hacking techniques, and open-source tools, CIPHER mimics expert reasoning and enhances the penetration testing process. It offers detailed explanations and guidance, particularly benefiting users with limited penetration testing experience. CIPHER is designed to be used specifically to guide beginner proceed with penetration testing a vulnerable machine for ethical research.
@@
Building on this foundation, we introduced the Findings, Action, Reasoning, and Results (FARR) Flow, a novel method for the automated evaluation of large language models in penetration testing scenarios. FARR Flow acts as a structured benchmark, simulating realistic pentesting tasks to rigorously assess the model’s technical knowledge, reasoning abilities, and practical effectiveness. By focusing on automated evaluation, FARR Flow provides a robust and realistic standard for measuring LLM performance in penetration testing, ensuring models are thoroughly tested in practical, real-world contexts.
@@
  • The methodology for CIPHER’s development, utilizing penetration testing write-ups structured as beginner- and expert-level conversations to train a large language model specifically for practical penetration testing assistance.
  • A novel augmentation technique called Findings, Action, Reasoning, and Results (FARR) Flow that condenses these write-ups into structured formats for more efficient evaluation.
  • The development and evaluation of an automated pentesting simulation benchmark based on FARR Flow, providing a rigorous and realistic framework for evaluating LLM performance in practical penetration testing guidance.
  • A comprehensive evaluation demonstrating CIPHER’s effectiveness and superior performance over existing LLMs in providing accurate guidance in FARR Flow and MITRE ATT&CK capabilities from PurpleLlama.
These contributions bridge the gap between AI advancements and practical cybersecurity applications, offering a new penetration testing domain-specific language model and reproducible benchmark for realistic penetration testing scenarios with consistent results.
@@ 2.1.1 Large Language Models
Transformers: Introduced by Vaswani et al. Vaswani et al. (2017) in “Attention is All You Need” (2017), Transformers have reshaped NLP by employing self-attention mechanisms to process entire sequences concurrently. They excel at capturing long-range dependencies, outperforming sequential models like RNNs. Transformers comprise self-attention layers, feed-forward networks, and positional encoding to maintain word order. Variants such as BERT and GPT have set benchmarks in NLP and extended into computer vision tasks through Vision Transformers (ViTs).
Open-Source LLMs: Open-source LLMs like the Llama family by Meta AI (Toronto, Canada) Touvron et al. (2023), Mistral (Paris, France) Jiang et al. (2023), Yi 1.5 by BAAI (Beijing, China) 01.AI et al. (2024), and Qwen 1.5 by Alibaba Group (Hangzhou, China) Bai et al. (2023) democratize access to powerful language models. Evolving from Llama 1 to Llama 3, these models increase token size and context length to enhance language understanding. Innovations like the mixture of experts in Mixtral by Mistral (Paris, France) Jiang et al. (2024) and unique training methods in Yi 1.5 push the boundaries further. Qwen 1.5 excels in multilingual contexts and incorporates agentic frameworks Qwen Team (2024), releasing smaller and powerful models qwe , making them versatile tools Touvron et al. (2023); Mann et al. (2020).
@@ 2.1.2 Applications and Challenges in Cybersecurity
In cybersecurity, large language models (LLMs) such as GPT-4 Achiam et al. (2023), PentestGPT Deng et al. (2023), hackingBuddyGPT Happe and Cito (2023), and HackerGPT Metta et al. (2024) have been deployed to assist with penetration testing and other security tasks. These models can analyze and synthesize large volumes of information, aiding in identifying vulnerabilities. However, using such advanced models poses significant challenges. The high cost and privacy risks associated with proprietary models like GPT-4, which handle sensitive vulnerability data, are significant concerns. Moreover, general-purpose LLMs often lack the specialized knowledge required for effective penetration testing.
@@ 2.2 Existing Tools and Methodologies for Penetration Testing
2.2.1 Manual and Automated Tools
Manual tools such as Netcat, Wireshark, and Burpsuite Community Edition offer robust vulnerability discovery and exploitation frameworks. These tools, however, heavily rely on user expertise to interpret results and often require manual intervention to execute complex attack scenarios.
Meanwhile, automated tools have emerged to streamline the penetration testing process by automating vulnerability scanning and initial exploitation attempts. Examples including Metasploit Holik et al. (2014), OpenVAS Aksu et al. (2019), and Nessus Kumar (2014) use predefined algorithms to detect and exploit common vulnerabilities. While these tools enhance efficiency, they may overlook subtle vulnerabilities that require human intuition and context to identify.
Combining manual and automated tools optimizes penetration testing effectiveness. Manual methods leverage human expertise to craft targeted attacks and interpret results, while automation accelerates routine tasks and broadens the scope of vulnerability detection. This hybrid approach ensures comprehensive security assessments addressing known vulnerabilities and emerging threats.
@@ 2.2.2 Advancements in AI for Penetration Testing
Recent advancements have seen the development of AI-driven tools designed to enhance penetration testing. Research efforts like PentestGPT Deng et al. (2023), AutoAttacker Xu et al. (2024), hackingBuddyGPT Happe and Cito (2023), ReaperAI Valencia (2024), PenHeal Huang and Zhu (2024), Pentest Copilot Goyal et al. (2024), and BreachSeek Alshehri et al. (2024) leverage LLMs for various penetration testing tasks. Despite their potential, these tools face several limitations:
@@
  • Inaccurate Operations and Commands: LLMs may generate inaccurate commands or misinterpret outputs, leading to ineffective testing Fang et al. (2024a, b).
@@ 2.2.3 Leveraging LLMs for Penetration Testing
On the other hand, general LLMs Achiam et al. (2023) used as penetration testing agents Deng et al. (2023); Happe and Cito (2023); Fang et al. (2024a, b) provide greater versatility and scalability. These models can perform a broader range of tasks beyond penetration testing and can be scaled easily by integrating additional modules or upgrading to more advanced general models. However, they often require complex setups involving various frameworks, modules, and pipelines, which can be resource-intensive and challenging to manage. The increased complexity can introduce vulnerabilities and make troubleshooting more difficult. Without significant adjustments and configurations, general LLMs may not achieve the same performance and specialization as native penetration testing models.
Table 1: Comparison of current penetration testing LLMs research. Notes (: available; : not available).
Name Native LLM Chat Assistant Foothold Priv. Escalation Multiple Tools
LLM as Hackers Happe et al. (2024)
@@
hackingBuddyGPT Happe and Cito (2023)
ReaperAI Valencia (2024)
PenHeal Huang and Zhu (2024)
Pentest Copilot Goyal et al. (2024)
BreachSeek Alshehri et al. (2024)
LLM Agents Can Autonomously Fang et al. (2024a, b)
@@
Given these considerations, we develop CIPHER as a native LLM tailored for penetration testing chatbot assistant. This approach ensures our model is optimized for the unique requirements of penetration testing tasks, thereby achieving higher accuracy and efficiency. A specialized model integrates domain-specific knowledge and tools more effectively, simplifying deployment and reducing the need for extensive customization. Additionally, handling sensitive data locally enhances security and minimizes the risk of exposure to vulnerable information.
@@ 2.3.2 Evaluation and Benchmarking Challenges
Current evaluation frameworks for cybersecurity models, such as PurpleLlama CyberSecEval Bhatt et al. (2024) with MITRE ATT&CK, primarily focus on predefined scenarios and known attack patterns Strom et al. (2018). These do not adequately measure a model’s ability to discover novel vulnerabilities or adapt to dynamic situations encountered in real-world penetration testing. Previous works like PentestGPT and AutoAttacker lack reproducible benchmarks, relying on manual testing and human scoring, which introduces bias and limits reproducibility Deng et al. (2023); Xu et al. (2024). Developing standardized, automated benchmarks that reflect real-world conditions remains a significant challenge for advancing the evaluation of cybersecurity LLMs.
@@ 2.3.3 Reproducible Benchmark for Penetration Testing Scenario
To address these challenges, we introduce the FARR Flow augmentation method, which automates the generation of penetration testing scenarios using high-quality write-ups. Unlike other existing penetration testing benchmarks for LLM, FARR Flow evaluation covers a wide range of tool usage capabilities, from reconnaissance to exploitation tools. FARR Flow evaluation results are also easily reproducible due to the scoring automation by using judge LLM. We also release our code on GitHub. This approach provides a dynamic and reproducible benchmark, assessing the LLM’s ability to reason and act like an expert penetration tester.
As shown in Table 2, our research demonstrates that FARR Flow evaluation fills the existing gap in current penetration testing benchmark for an LLM that is easily reproducible, open-source, and covers a wide range of tools.
Table 2: Comparison of penetration testing evaluation approaches. Notes (: yes; : not provided).
@@
hackingBuddyGPT Happe and Cito (2023)
ReaperAI Valencia (2024)
PenHeal Huang and Zhu (2024)
Pentest Copilot Goyal et al. (2024)
BreachSeek Alshehri et al. (2024)
LLM Can Autonomously Fang et al. (2024a, b)
@@
In Section 4.3, we propose FARR Flow augmentation and evaluation, which utilize existing high-quality penetration testing write-ups as a base for penetration testing flow. Therefore, creating the correct step of penetration testing of a specific vulnerable machine is no longer needed, enabling us to measure the performance of the model on existing vulnerable machines. Unlike traditional human scoring, which cannot be consistently reproducible, in FARR Flow we can change the evaluator model to avoid and minimize bias using multiple SOTA models as evaluators.